X^2-4x+2=0
The x^2 means x squared

Answer by krutarthas(22)

x²-4x+2=0
⇒x²-4x+2+2=0+2
⇒x²-2.x.2+2²=2
⇒(x-2)²=2
⇒x-2=√2
⇒x=√2+2